# Ego vs. Exo and Active vs. Passive: Investigating the Effects of Viewpoint and Navigation on Spatial Immersion and Understanding in Immersive Storytelling

### v1

@@ Abstract.
Four Combinations of Viewpoints and Navigations. The first scene (leftmost) depicts the person standing in a classroom setting with several figures seated around tables. The figures appear to be rendered in a basic gray, suggesting that they are part of a virtual environment. The second scene shows the person standing outside of what appears to be a virtual building, interacting with a door or window. The environment has a simplistic, grayish brick design. The third scene presents the person holding a 3D model of the Earth, as though they are interacting with a virtual globe. The final scene (rightmost) shows the person seemingly holding or interacting with a floating model of a cityscape, with buildings arranged on a platform that hovers in the virtual space.
@@ 1. Introduction
Visual storytelling uses the combination of visual elements and narratives to communicate complex ideas (e.g., data, facts, and opinions) in a way that is more engaging and descriptive than traditional text-only stories (Kulkarni et al., 2023). This approach has been widely used in various fields, including journalism (Times, 2024c; Post, 2024; Guardian, 2024), education (Williams, 2019; Baharuddin et al., 2023), commercial activities (Moin et al., 2020; Zhou et al., 2005), and scientific communication (Ma et al., 2012). It plays a crucial role in enhancing the readers’ understanding of content, as well as increasing their emotional engagement and connection to the story.
@@
The evolution of digital technologies has greatly expanded the possibilities of visual storytelling, offering new genres and more diverse ways to deliver content (Segel and Heer, 2010). The introduction of digital displays brought dynamic and interactive visual content, changing how readers interact and engage with stories (Yu and Shi, 2018; Greussing et al., 2020). The rise of mobile devices further transformed access and interaction modalities, enabling readers to experience stories anywhere with new input modalities such as touch-based interactions (Lima et al., 2020; Sheremetieva et al., 2022). As digital technologies continue to evolve, the pressing question now is: ”What is next for visual storytelling, and how can it offer novel experiences?”
@@
As immersive technologies rapidly develop, they are becoming increasingly accessible and affordable, positioning them as promising candidates for the next generation of major public digital platforms. Immersive devices (e.g., Virtual Reality (VR) headsets) can render 2D and 3D information in the surrounding space or even replace the users’ entire visual field, offering unprecedented opportunities for visual storytelling (Kraus et al., 2021; Isenberg et al., 2018). A prominent use of immersive technology for storytelling is 360-degree videos (Google, 2016; Illustrated, 2017). However, these videos often limit the readers’ navigation, providing little control over story progression or detail exploration due to their fully author-driven nature. Beyond static videos, major news outlets like The New York Times and The Washington Post have embraced interactive formats in their web-based articles to present public events (e.g., COVID-19 and elections) more dynamically. They also leverage immersive technologies for more interactive experiences, such as mobile AR stories created by The New York Times (Times, 2022). However, these narratives tend to rely on basic touchscreen inputs and offer limited immersion. They are often confined to a single scene without transitions between multiple narrative elements. In this work, we focus on improving immersive storytelling specifically for public journalism. We seek to extend the design of these journalistic stories into immersive environments, where interactivity and complete storylines can foster deeper engagement and understanding.
@@
Our goal is to leverage the unique display and interaction affordances of immersive technologies in storytelling. A major advantage of these technologies is their ability to render 3D spatial information at its original scale (Isenberg et al., 2018). Building on this characteristic, we investigated how people visually perceive 3D spatial representations (i.e., ego- vs. exocentric viewpoints) and how they can interactively navigate in the 3D space (i.e., active vs. passive navigation) as an initial exploration. Specifically, readers can experience an immersive story from an exocentric (exo) viewpoint to gain an overview of a 3D scene or from an egocentric (ego) viewpoint to fully immerse themselves with life-sized artifacts. Additionally, readers can either actively move in space and interact with digital artifacts with greater embodiment and agency or passively follow predefined transitions. We anticipate that these two factors (i.e., viewpoint and navigation) will have a significant effect on user-perceived spatial immersion and understanding in immersive storytelling. Our research goal is to understand the advantages and limitations of these design choices in spatial immersion and understanding in immersive VR stories.
@@
To study the effects of viewpoints and navigation in immersive storytelling, we needed to develop immersive stories that were controlled for these factors. While there are many web-based stories for public journalism, immersive versions are still rare. Since we aimed to focus on viewpoint and navigation rather than the story content itself, we leveraged existing web-based stories and created versions in immersive environments to cover the design variations we needed for the study. We collected web-based stories from major news outlets, particularly those with rich 3D spatial content, such as a New York Times story about how COVID-19 spreads in a classroom (Bartzokas et al., 2021). These visual stories are in single scene and follow a linear structure. We consider this atomicity as a starting point for studying our intended effects. Meanwhile, this basic structure serves as the foundational unit for more complex stories. We then conducted a formative study with three VR experts to elicit potential designs of 117 selected stories in an immersive format, with an intended focus on viewpoint and navigation. Based on the design considerations from this formative study, we adapted four representative stories into immersive formats. Finally, we conducted a user study with 24 participants to compare the 2 combinations (ego vs. exo viewpoint, and active vs. passive navigation). Our paper structure follows the steps we took to prepare and conduct the study, as illustrated in Figure 2.
A research workflow diagram illustrating the process of a VR storytelling study. The left side details the Formative Study (Section 3), starting with web-based story collection and VR design elicitation from 3 experts to identify design considerations. These considerations are applied to the VR Story Adaptation (Section 4), involving design and implementation of immersive stories. The right side describes the User Study (Section 5), which evaluates different combinations of viewpoint (egocentric vs. exocentric) and navigation (active vs. passive) with 24 participants.
@@ 2. Related Works
Visual Storytelling. Visual storytelling is widely used in journalism to help audiences understand data-backed facts through static or dynamic visual elements, such as images, videos, and animations. Segal and Heer (Segel and Heer, 2010) were among the first to summarize emerging narrative visualization techniques in web-based storytelling, discussing design variations across different genres. Hullman et al. (Hullman and Diakopoulos, 2011) studied how information prioritization affects user interpretation. Kosara and Mackinlay (Kosara and Mackinlay, 2013) further highlighted the potential of storytelling in visualizations, emphasizing its promising future. Subsequent research explored design considerations in visual storytelling, proposing guidelines to improve content representation and interaction. For instance, Morais et al.(Morais et al., 2020) proposed a design framework for anthropographic data stories, while Dasu et al. (Dasu et al., 2023) examined the use of characters in data stories to enhance enjoyment and persuasion. Yang et al. (Yang et al., 2021) also explored the application of narrative structures in data stories. These studies connect design practices with human cognition theories, offering valuable insights for visual storytelling design.
@@
Beyond narrative structure, audience engagement and interaction are critical to the story’s impact, persuasion, and memorability. With advancements in technology, visual storytelling has expanded to include videos, games, and VR/AR platforms (Zhao and Elmqvist, 2023; Méndez et al., 2025; Zhou et al., 2023), offering audiences more interactive experiences than static texts and images alone. VR/AR, in particular, immerses users in virtual spaces, enhancing spatial coherence and unlocking new design possibilities. This paper focuses on two key elements in VR/AR and visual storytelling: viewpoints and navigation, which shape user perspective and agency, directly influencing immersion, comprehension, and emotional engagement. We aim to identify effective design practices for integrating these aspects into immersive storytelling.
@@
Immersive Experiences. The goals of visual storytelling are multifaceted (Amini et al., 2018). For authors, common objectives include (a) communicating facts and insights, and (b) persuading audiences of the opinions conveyed through the story. For audiences, the primary goal is to gain new knowledge, such as news, insights, and perspectives, while entertainment also plays a significant role, as stories may be consumed for enjoyment.
@@
While 2D screens are effective for visual storytelling, immersive technologies like VR/AR extend stories into 3D, closing the gap between audiences and narratives by placing them in the same space. Immersive experiences offer several advantages. First, they enhance presence and immersion (Servotte et al., 2020). VR/AR can simulate real-world scenarios (Zhao et al., 2023) with real-time rendering and embodied interactions (Huang et al., 2023; Zhu et al., 2024a; In et al., 2023), or even go beyond reality (e.g., ground-level scaling (Abtahi et al., 2019)). Second, they improve understanding and retention (Motejlek and Alpay, 2023; Yildirim et al., 2019). VR/AR is frequently used in scientific visualizations (Marriott et al., 2018), helping users extract and communicate insights more effectively. Third, they enhance emotional communication (Kandaurova and Lee, 2019), narrowing the emotional distance between authors and audiences and fostering resonance and reflection on the story’s themes (Bahng et al., 2020).
@@
Immersive Storytelling. Immersive storytelling, a broader topic of integrating immersive experiences to different storytellings, has been studied in various other contexts. In AR, Shin et al. (Shin et al., 2019) studied the two factors: room size and furniture density on a AR crime-solving game, and they suggested more context-awareness such as using virtual objects as substitutes in large and empty rooms. A later survey (Shin et al., 2024) furthered this idea, proposing to balance virtual and real experiences through situated user interactions. In VR, Lee et al. leveraged VR to materialize abstract measure and units so that audiences could connect with their natural experiences (Lee et al., 2020). Hall et al. focused on synchronous immersive data presentation and communication (Hall et al., 2022). VR storytelling is particularly relevant in cinematic. Collen et al. investigated the cinematic techniques used in immersive narrative visualizations of 3D spatial information (Conlen et al., 2023). Practically, Wu and Karwas (Wu and Karwas, 2024) developed a technique that allows free viewpoint change and with time-based interactivity. Other immersive storytelling works focusing on specific applications, such as situated awareness of social problems (Assor et al., 2024; Zhu et al., 2024b) or emotional enjoyment of visualizations (Romat et al., 2020). These works contribute to the design of immersive storytelling by highlighting the benefits of immersive technologies we should leverage in specific topics or contexts, but none explored the representation and interaction design on visual jouralism (e.g. scrollytelling) in immersive environment.
@@
There are some studies focusing on how traditional storytellings on 2D screens can be transported to VR/AR, which are primarily based on the video format (Yang et al., 2023; Zollmann et al., 2020; Nash, 2018). The adaptation of interactive visual stories, a popular form of visual storytelling, are still underexplored.
@@
Viewpoint and Navigation in Immersive Environments. In traditional digital media (e.g., computer monitors or mobile devices), content is typically viewed from an exocentric (third-person) viewpoint, offering a comprehensive overview of spatial relationships and layout. Immersive technology, while capable of maintaining an exocentric view, uniquely allows for a 360-degree stereoscopic experience from an egocentric (first-person) viewpoint. Despite this capability, the egocentric view is not widely adopted due to mixed findings on its effectiveness. Yang et al. (Yang et al., 2018) found that exocentric views significantly outperformed egocentric views for geographic analysis tasks. Similarly, Kraus et al. (Kraus et al., 2019) and Yang et al. (Yang et al., 2020a) noted limitations in egocentric 3D scatterplot visualizations, such as lack of overview and targets being out of view. However, the egocentric view has shown promise in multi-view management within immersive environments (In et al., 2024; Davidson et al., 2022; Satriadi et al., 2020; Luo et al., 2022), and studies have demonstrated improved memory and knowledge retention in egocentric settings (Krokos et al., 2019; Yang et al., 2020b). Notably, Hoppe et al. (Hoppe et al., 2022) argued for a perspective continuum in immersive environments, finding no difference in users’ perceived workload, sense of presence, and engagement between egocentric and exocentric viewpoints in a VR combat game they developed. However, it remains questionable whether this perspective continuum is applicable in immersive storytelling, as storytelling involves different perception and cognitive processes compared to games. Given these mixed results, our research explores the advantages and drawbacks of both exocentric and egocentric viewpoints in the context of visual storytelling.
@@
Navigation in immersive environments refers to the process of moving to the target location within 3D virtual environments. It is a crucial interaction for exploring 3D spaces in these settings. In the context of immersive storytelling, navigation plays a vital role in allowing users to visually explore different parts of the story and advance the narrative (e.g., a story progresses to the next stage once the viewer reaches a specific location). Significant efforts have been devoted to developing various navigation techniques, including gesture design (Tursunov et al., 2024), utilizing different input modalities (Świdrak et al., 2024), leveraging perception and space manipulation (Dong et al., 2021), and experimenting with different scales (Mirhosseini et al., 2019; Abtahi et al., 2019). Luca et al. (Di Luca et al., 2021) conducted an extensive survey of these techniques and compiled a Locomotion Vault. However, some of these innovative techniques require additional hardware, while others demand a steep learning curve. As a result, the mainstream and default navigation techniques adopted by major platforms remain natural walking and teleportation. Considering the accessibility aspect of storytelling, we decided to focus on using mainstream navigation techniques for a broader user group. In many immersive applications, users need to actively navigate the 3D space to complete tasks. However, in storytelling, we have the option to let users passively follow a predefined path without explicit navigation actions (such as walking or teleportation). While active navigation provides more interactivity, it carries the risk of disorienting the user. In contrast, passive navigation maintains narrative control but may reduce the sense of agency and induce motion sickness. Deciding which method to use is a unique and important consideration for immersive storytelling. In relation to this topic, Lages and Bowman (Lages and Bowman, 2018) compared physical walking (where the user moves themselves) with grab-and-move techniques (where the user moves the view) in VR, demonstrating that performance may depend on the user’s spatial abilities. Although both methods they studied were active, we are particularly interested in the performance of passive navigation when combined with a predefined storyline.
@@ 3. Eliciting Design Considerations for Immersive Story Adaptations
A process flow diagram showing the story filtering process for a VR study. The flow begins with over 1,000 visual stories, from which static infographics are filtered out. A total of 148 dynamic visual stories are selected through Initial Story Collection. These dynamic stories are further filtered, excluding stories lacking sufficient 3D spatial information, resulting in 117 VR adaptable stories. Finally, through Elicitation, four design considerations (DCs) are identified: Main Visual Elements, Story Narratives, Progression Triggers, and Text and Visual Guidance.
@@
Visual storytelling can be characterized by four dimensions: narratives, transitions, structure, and the balance between explorability and explanability (Stolper et al., 2018; Thudt et al., 2018). With various options in each category, visual storytelling can become quite complex. Since immersive storytelling is an emerging field with limited publicly available resources and few established design guidelines, we decided to develop our immersive stories. Creating high-quality stories is a complex and challenging task. Therefore, instead of creating stories from scratch, we chose to adapt web-based stories from major news outlets. This approach allows us to ensure story quality while reducing potential confounding factors. Considering the aforementioned four dimensions, we found that most stories from major news outlets follow a single-scene linear narrative structure with simple and continuous transitions and predominantly feature author-driven explanability. This simplicity is ideal for our initial exploration into immersive storytelling. We further focused on stories with significant 3D spatial content due to their greater synergy with immersive environments. This approach enabled us to concentrate on the study itself, maintaining story quality while avoiding potential design biases.
@@
Initial story collection. Our first step is to collect abundant web-based stories. To ensure the quality of our visual story dataset, we collected stories from major news outlets and popular online media platforms, including The New York Times (NYT) (Times, 2024c), The Washington Post (Post, 2024), Bloomberg (Bloomberg, 2024), The Guardian (Guardian, 2024), Reuters (Reuters, 2024), National Geographic (Geographics, 2024), Financial Times (Times, 2024a), Los Angeles Times (Times, 2024b) and The Pudding (Pudding, 2024). Some platforms feature dedicated visual story sections, which we prioritized, such as the Graphics section in NYT (Times, 2024c) and the Visual Stories section in The Washington Post (Post, 2024).
@@
A diagram illustrating key design considerations for immersive storytelling in VR, presented in four categories: 1) Main Visual Elements: Real-world transform (elements positioned based on real-world coordinates) vs. Exo transform (global context positioning). 2) Story Narrative Placement: World-fixed (narrative anchored in the virtual world), Floating panel (text on a floating interface), and Situated (contextual narrative placement). 3) Navigation Triggers: Locomotion (natural walking or teleportation) and Physical interactions (interacting with objects like opening a window). 4) Text and Visual Guidance: Text placement (World-fixed, Floating panel, or Situated) and Visual placement (User-oriented vs. Situated for guiding attention).

### v2

@@ 1. Introduction
Visual storytelling uses the combination of visual elements and narratives to communicate complex ideas (e.g., data, facts, and opinions) in a way that is more engaging and descriptive than traditional text-only stories (kulkarni2023innovating). This approach has been widely used in various fields, including journalism (nyt; wapo; guardian), education (Williams2019Attending; Baharuddin2023The), commercial activities (Moin2020Storytelling; Zhou2005VISUAL), and scientific communication (Ma2012Scientific). It plays a crucial role in enhancing the readers’ understanding of content, as well as increasing their emotional engagement and connection to the story.
@@
The evolution of digital technologies has greatly expanded the possibilities of visual storytelling, offering new genres and more diverse ways to deliver content (segel2010narrative). The introduction of digital displays brought dynamic and interactive visual content, changing how readers interact and engage with stories (Yu2018A; Greussing2020Learning). The rise of mobile devices further transformed access and interaction modalities, enabling readers to experience stories anywhere with new input modalities such as touch-based interactions (Lima2020Sketch-Based; Sheremetieva2022Touch). As digital technologies continue to evolve, the pressing question now is: ”What is next for visual storytelling, and how can it offer novel experiences?”
@@
As immersive technologies rapidly develop, they are becoming increasingly accessible and affordable, positioning them as promising candidates for the next generation of major public digital platforms. Immersive devices (e.g., Virtual Reality (VR) headsets) can render 2D and 3D information in the surrounding space or even replace the users’ entire visual field, offering unprecedented opportunities for visual storytelling (kraus2021value; isenberg2018immersive). A prominent use of immersive technology for storytelling is 360-degree videos (google_2016_beyond; sportsillustrated_2017_chapter). However, these videos often limit the readers’ navigation, providing little control over story progression or detail exploration due to their fully author-driven nature. Beyond static videos, major news outlets like The New York Times and The Washington Post have embraced interactive formats in their web-based articles to present public events (e.g., COVID-19 and elections) more dynamically. They also leverage immersive technologies for more interactive experiences, such as mobile AR stories created by The New York Times (thenewyorktimes_2022_showcasing). However, these narratives tend to rely on basic touchscreen inputs and offer limited immersion. They are often confined to a single scene without transitions between multiple narrative elements. In this work, we focus on improving immersive storytelling specifically for public journalism. We seek to extend the design of these journalistic stories into immersive environments, where interactivity and complete storylines can foster deeper engagement and understanding.
@@
Our goal is to leverage the unique display and interaction affordances of immersive technologies in storytelling. A major advantage of these technologies is their ability to render 3D spatial information at its original scale (isenberg2018immersive). Building on this characteristic, we investigated how people visually perceive 3D spatial representations (i.e., ego- vs. exocentric viewpoints) and how they can interactively navigate in the 3D space (i.e., active vs. passive navigation) as an initial exploration. Specifically, readers can experience an immersive story from an exocentric (exo) viewpoint to gain an overview of a 3D scene or from an egocentric (ego) viewpoint to fully immerse themselves with life-sized artifacts. Additionally, readers can either actively move in space and interact with digital artifacts with greater embodiment and agency or passively follow predefined transitions. We anticipate that these two factors (i.e., viewpoint and navigation) will have a significant effect on user-perceived spatial immersion and understanding in immersive storytelling. Our research goal is to understand the advantages and limitations of these design choices in spatial immersion and understanding in immersive VR stories.
@@
To study the effects of viewpoints and navigation in immersive storytelling, we needed to develop immersive stories that were controlled for these factors. While there are many web-based stories for public journalism, immersive versions are still rare. Since we aimed to focus on viewpoint and navigation rather than the story content itself, we leveraged existing web-based stories and created versions in immersive environments to cover the design variations we needed for the study. We collected web-based stories from major news outlets, particularly those with rich 3D spatial content, such as a New York Times story about how COVID-19 spreads in a classroom (covid). These visual stories are in single scene and follow a linear structure. We consider this atomicity as a starting point for studying our intended effects. Meanwhile, this basic structure serves as the foundational unit for more complex stories. We then conducted a formative study with three VR experts to elicit potential designs of 117 selected stories in an immersive format, with an intended focus on viewpoint and navigation. Based on the design considerations from this formative study, we adapted four representative stories into immersive formats. Finally, we conducted a user study with 24 participants to compare the 2 combinations (ego vs. exo viewpoint, and active vs. passive navigation). Our paper structure follows the steps we took to prepare and conduct the study, as illustrated in Figure 2.
@@ 2. Related Works
Visual Storytelling. Visual storytelling is widely used in journalism to help audiences understand data-backed facts through static or dynamic visual elements, such as images, videos, and animations. Segal and Heer (segel2010narrative) were among the first to summarize emerging narrative visualization techniques in web-based storytelling, discussing design variations across different genres. Hullman et al. (hullman2011visualization) studied how information prioritization affects user interpretation. Kosara and Mackinlay (kosara2013storytelling) further highlighted the potential of storytelling in visualizations, emphasizing its promising future. Subsequent research explored design considerations in visual storytelling, proposing guidelines to improve content representation and interaction. For instance, Morais et al.(morais2020showing) proposed a design framework for anthropographic data stories, while Dasu et al. (dasu2023character) examined the use of characters in data stories to enhance enjoyment and persuasion. Yang et al. (yang2021design) also explored the application of narrative structures in data stories. These studies connect design practices with human cognition theories, offering valuable insights for visual storytelling design.
@@
Beyond narrative structure, audience engagement and interaction are critical to the story’s impact, persuasion, and memorability. With advancements in technology, visual storytelling has expanded to include videos, games, and VR/AR platforms (zhao2023stories; mendez2025immersive; zhou2023data), offering audiences more interactive experiences than static texts and images alone. VR/AR, in particular, immerses users in virtual spaces, enhancing spatial coherence and unlocking new design possibilities. This paper focuses on two key elements in VR/AR and visual storytelling: viewpoints and navigation, which shape user perspective and agency, directly influencing immersion, comprehension, and emotional engagement. We aim to identify effective design practices for integrating these aspects into immersive storytelling.
@@
Immersive Experiences. The goals of visual storytelling are multifaceted (amini2018evaluating). For authors, common objectives include (a) communicating facts and insights, and (b) persuading audiences of the opinions conveyed through the story. For audiences, the primary goal is to gain new knowledge, such as news, insights, and perspectives, while entertainment also plays a significant role, as stories may be consumed for enjoyment.
@@
While 2D screens are effective for visual storytelling, immersive technologies like VR/AR extend stories into 3D, closing the gap between audiences and narratives by placing them in the same space. Immersive experiences offer several advantages. First, they enhance presence and immersion (servotte2020vr). VR/AR can simulate real-world scenarios (zhao2023leanon) with real-time rendering and embodied interactions (huang2023embodied; zhu2024compositingvis; in2023table), or even go beyond reality (e.g., ground-level scaling (abtahi2019m)). Second, they improve understanding and retention (Motejlek2023The; Yildirim2019The). VR/AR is frequently used in scientific visualizations (marriott2018immersive), helping users extract and communicate insights more effectively. Third, they enhance emotional communication (kandaurova2019effects), narrowing the emotional distance between authors and audiences and fostering resonance and reflection on the story’s themes (bahng2020reflexive).
@@
Immersive Storytelling. Immersive storytelling, a broader topic of integrating immersive experiences to different storytellings, has been studied in various other contexts. In AR, Shin et al. (shin2019any) studied the two factors: room size and furniture density on a AR crime-solving game, and they suggested more context-awareness such as using virtual objects as substitutes in large and empty rooms. A later survey (shin2024investigating) furthered this idea, proposing to balance virtual and real experiences through situated user interactions. In VR, Lee et al. leveraged VR to materialize abstract measure and units so that audiences could connect with their natural experiences (lee2020data). Hall et al. focused on synchronous immersive data presentation and communication (hall2022augmented). VR storytelling is particularly relevant in cinematic. Collen et al. investigated the cinematic techniques used in immersive narrative visualizations of 3D spatial information (conlen2023cinematic). Practically, Wu and Karwas (wu2024metamorphosis) developed a technique that allows free viewpoint change and with time-based interactivity. Other immersive storytelling works focusing on specific applications, such as situated awareness of social problems (assor2024augmented; zhu2024reader) or emotional enjoyment of visualizations (romat2020dear). These works contribute to the design of immersive storytelling by highlighting the benefits of immersive technologies we should leverage in specific topics or contexts, but none explored the representation and interaction design on visual jouralism (e.g. scrollytelling) in immersive environment.
@@
There are some studies focusing on how traditional storytellings on 2D screens can be transported to VR/AR, which are primarily based on the video format (yang2023understanding; Zollmann2020CasualVRVideos; Nash2018Virtually). The adaptation of interactive visual stories, a popular form of visual storytelling, are still underexplored.
@@
Viewpoint and Navigation in Immersive Environments. In traditional digital media (e.g., computer monitors or mobile devices), content is typically viewed from an exocentric (third-person) viewpoint, offering a comprehensive overview of spatial relationships and layout. Immersive technology, while capable of maintaining an exocentric view, uniquely allows for a 360-degree stereoscopic experience from an egocentric (first-person) viewpoint. Despite this capability, the egocentric view is not widely adopted due to mixed findings on its effectiveness. Yang et al. (yang2018maps) found that exocentric views significantly outperformed egocentric views for geographic analysis tasks. Similarly, Kraus et al. (kraus2019impact) and Yang et al. (yang2020embodied) noted limitations in egocentric 3D scatterplot visualizations, such as lack of overview and targets being out of view. However, the egocentric view has shown promise in multi-view management within immersive environments (in2024evaluating; davidson2022exploring; satriadi2020maps; luo2022should), and studies have demonstrated improved memory and knowledge retention in egocentric settings (krokos2019virtual; yang2020virtual). Notably, Hoppe et al. (hoppe2022there) argued for a perspective continuum in immersive environments, finding no difference in users’ perceived workload, sense of presence, and engagement between egocentric and exocentric viewpoints in a VR combat game they developed. However, it remains questionable whether this perspective continuum is applicable in immersive storytelling, as storytelling involves different perception and cognitive processes compared to games. Given these mixed results, our research explores the advantages and drawbacks of both exocentric and egocentric viewpoints in the context of visual storytelling.
@@
Navigation in immersive environments refers to the process of moving to the target location within 3D virtual environments. It is a crucial interaction for exploring 3D spaces in these settings. In the context of immersive storytelling, navigation plays a vital role in allowing users to visually explore different parts of the story and advance the narrative (e.g., a story progresses to the next stage once the viewer reaches a specific location). Significant efforts have been devoted to developing various navigation techniques, including gesture design (tursunov2024creating), utilizing different input modalities (swidrak2024beyond), leveraging perception and space manipulation (dong2021tailored), and experimenting with different scales (mirhosseini2019exploration; abtahi2019m). Luca et al. (di2021locomotion) conducted an extensive survey of these techniques and compiled a Locomotion Vault. However, some of these innovative techniques require additional hardware, while others demand a steep learning curve. As a result, the mainstream and default navigation techniques adopted by major platforms remain natural walking and teleportation. Considering the accessibility aspect of storytelling, we decided to focus on using mainstream navigation techniques for a broader user group. In many immersive applications, users need to actively navigate the 3D space to complete tasks. However, in storytelling, we have the option to let users passively follow a predefined path without explicit navigation actions (such as walking or teleportation). While active navigation provides more interactivity, it carries the risk of disorienting the user. In contrast, passive navigation maintains narrative control but may reduce the sense of agency and induce motion sickness. Deciding which method to use is a unique and important consideration for immersive storytelling. In relation to this topic, Lages and Bowman (lages2018move) compared physical walking (where the user moves themselves) with grab-and-move techniques (where the user moves the view) in VR, demonstrating that performance may depend on the user’s spatial abilities. Although both methods they studied were active, we are particularly interested in the performance of passive navigation when combined with a predefined storyline.
@@ 3. Eliciting Design Considerations for Immersive Story Adaptations
Visual storytelling can be characterized by four dimensions: narratives, transitions, structure, and the balance between explorability and explanability (stolper2018data; thudt2018exploration). With various options in each category, visual storytelling can become quite complex. Since immersive storytelling is an emerging field with limited publicly available resources and few established design guidelines, we decided to develop our immersive stories. Creating high-quality stories is a complex and challenging task. Therefore, instead of creating stories from scratch, we chose to adapt web-based stories from major news outlets. This approach allows us to ensure story quality while reducing potential confounding factors. Considering the aforementioned four dimensions, we found that most stories from major news outlets follow a single-scene linear narrative structure with simple and continuous transitions and predominantly feature author-driven explanability. This simplicity is ideal for our initial exploration into immersive storytelling. We further focused on stories with significant 3D spatial content due to their greater synergy with immersive environments. This approach enabled us to concentrate on the study itself, maintaining story quality while avoiding potential design biases.
@@
Initial story collection. Our first step is to collect abundant web-based stories. To ensure the quality of our visual story dataset, we collected stories from major news outlets and popular online media platforms, including The New York Times (NYT) (nyt), The Washington Post (wapo), Bloomberg (bloomberg), The Guardian (guardian), Reuters (reuters), National Geographic (NG), Financial Times (FT), Los Angeles Times (lat) and The Pudding (pudding). Some platforms feature dedicated visual story sections, which we prioritized, such as the Graphics section in NYT (nyt) and the Visual Stories section in The Washington Post (wapo).